OpenStudy (anonymous):

how do I convert 89 to a decimal?

OpenStudy (anonymous):

0.89

OpenStudy (saifoo.khan):

how many decimal places?

OpenStudy (anonymous):

2 this book says it's 9.434 but I do not understand how this is true

OpenStudy (saifoo.khan):

u must write the whole ques here!

OpenStudy (anonymous):

It's a story prob. based on Pythagorean theorem

OpenStudy (anonymous):

The triangle is 8ft on one side and 5ft on the bottom

OpenStudy (saifoo.khan):

can u post the photo over here? take a picture and upload it here

OpenStudy (anonymous):

I can sure try :) give me a min

OpenStudy (saifoo.khan):

kk, np!

OpenStudy (anonymous):

Ok. I can't figure it out. Here's the problem though. Darla is building a new desk. To make sure the had made a square corner, she measures 8ft from the corner along one edge and 5ft from the corner along the other edge. HOw long should the diagonal be between those tow points if the corner is a right angle?

OpenStudy (anonymous):

it also says c^2=89 but I'm suppose to find the decimal, which is 9.434. I just dont understand how they got the decimal

OpenStudy (saifoo.khan):

Yea its correct.

OpenStudy (saifoo.khan):

\[\Huge 8^2 + 5^2 = c^2 \]

OpenStudy (anonymous):

square root of 89

OpenStudy (saifoo.khan):

\[\Huge c^2 = 89\]

OpenStudy (saifoo.khan):

\[\Huge \sqrt{89}\]

OpenStudy (anonymous):

wow!! I feel so ignorant! lol

OpenStudy (anonymous):

I definitely see now! Thanks so much you guys!

OpenStudy (anonymous):

next time word it better ;)

OpenStudy (anonymous):

I know. sorry =)

OpenStudy (saifoo.khan):

Uw. :)